OpenStudy (anonymous):

solve the differential equation dy/dx=-csc^2(5x)/(cot(5x)) for the particular solution that passes through the point (pi/20,3)

OpenStudy (kc_kennylau):

\[\frac{dy}{dx}=-\frac{\csc^2(5x)}{\cot(5x)},y(\frac\pi{20})=3\]

OpenStudy (kc_kennylau):

Simplify the stuff by expressing everything in terms of sine and cosine first

OpenStudy (anonymous):

what would the first step be? im confused on how to start it

OpenStudy (kc_kennylau):

Simplify \(-\dfrac{\csc^2(5x)}{\cot(5x)}\) first

OpenStudy (anonymous):

i dont know how to simplify that :/

OpenStudy (kc_kennylau):

By expressing everything in terms of sine and cosine

OpenStudy (kc_kennylau):

\[\tan\theta=\frac{\sin\theta}{\cos\theta}\]\[\cot\theta=\frac{\cos\theta}{\sin\theta}\]\[\csc\theta=\frac1{\sin\theta}\]\[\sec\theta=\frac1{\cos\theta}\]

OpenStudy (unklerhaukus):

\[\frac{\csc^2(\theta)}{\cot(\theta)}=\frac{1/\sin^2(\theta)}{\cos(\theta)/\sin(\theta)}=\] simplify this by multiplying the numerator and denominator by sine theta

OpenStudy (kc_kennylau):

For example, \(\cot(5x)=\dfrac{\cos(5x)}{\sin(5x)}\)

OpenStudy (anonymous):

whats next?

OpenStudy (kc_kennylau):

Follow his step

OpenStudy (anonymous):

\[\frac{ 1 }{ \sin(5x) }\]/cos(5x)

OpenStudy (unklerhaukus):

Use this Product to Sum formula in the numerator, to get it into a single trig function\[\sin\psi\cos\phi=\tfrac12[\sin(\psi+\phi)+\sin(\psi-\phi)]\]

OpenStudy (unklerhaukus):

*denominator

OpenStudy (unklerhaukus):

.... @iTzVinny21